OpenStudy (anonymous):

a vehicle uses 1 tablespoon of gasoline to drive 125 yards. how many miles can the vehicle travel per gallon? round the answer to the nearest mile. (256 tablespoon= 1 gallon.)

OpenStudy (anonymous):

how many yards in a mile?

OpenStudy (anonymous):

1760

OpenStudy (anonymous):

18.18 miles

OpenStudy (anonymous):

use conversion factors such that the units "cancel out" in a numerator and denominator and the only remaining are miles/gallon \[\frac{ 125 yards }{ 1 tbsp } \times \frac{ 1760 miles }{ 1 yard } \times \frac{ 256 tbsp }{ 1 gallon} = \frac{X miles }{ 1 gallon }\]